The element lawrencium has a half-life of 10 hours. What fraction of 1 ounce of lawrencium will remain after 4 days 4 hours?
Answer by stanbon(75887)

The element lawrencium has a half-life of 10 hours. What fraction of 1 ounce of lawrencium will remain after 4 days 4 hours?
-----
4 days and 4 hrs = 4*24+4 = 100 hrs
-------
A(t) = Ao(1/2)^(t/10)
-------
A(100) = 1*(1/2)^(100/10) = (1/2)^10 = (1/1024)ounces
====================
Cheers,
Stan H.